The sum of three consecutive odd integers equals 240 more than the third integer. Find all three integers
Answer by pedjajov(51)

If first odd number is x, then second consecutive odd number is x+2 and third is x+4.
So we have

x+x+2+x+4 = x+4 + 240
2x+2=240
2x=238
x=119, this is first number

x+2=121, this is second number

x+4=123, this is third number

check:
119+121+123=363
123+240 =363, OK